### v2

# Toward a Deterministic Nucleation Theory for Chirality-Controlled Nanotube Synthesis

Abstract
The electronic properties of carbon nanotubes are dictated by their chirality, specified by the integer indices $(n,m)$. The absence of a predictive theory directly linking catalysts to the resulting $(n,m)$ has hindered the controlled synthesis of specific chiralities. Here we narrow this gap through a two-step approach. First, we derive a universal vector sum rule that quantifies the relationship between cap structure and chirality, demonstrating that $(n,m)$ is uniquely determined by the vector sum of the six pentagon coordinates within the cap. The complete set of these coordinates, evaluated across rotationally symmetric lattice frames, fully defines the carbon cap structure. Building on this topological foundation, we establish that chirality is encoded deterministically at nucleation, as perturbations to pentagon positions lead to predictable, quantifiable, and kinetically irreversible shifts in $(n,m)$. By examining existing experimental evidence, we further propose that the preferential formation of $(12,6)$ nanotubes arises from a six-fold symmetric cap with epitaxial matching to the $\langle 1,1,1\rangle$ catalyst facet. Our work provides a theoretical framework that redefines the field, shifting the paradigm from stochastic growth kinetics to deterministic nucleation programming and paving the way toward predictable synthesis.
Carbon nanotubes (CNTs) hold exceptional promise for semiconductor applications due to their extraordinary electrical conductivity and chirality-dependent band gaps [1, 2]. However, achieving chiral-selective synthesis remains a fundamental challenge in nanotechnology [3, 4, 5]. Chemical vapor deposition with solid catalysts offers a scalable and promising approach toward chirality-specific growth, having produced several CNTs with remarkable enantiomeric purity—including $(12,6)$ at 92% and $(14,4)$ at 99% [6, 7]. Despite these advances, the physical principles governing chiral preference remain poorly understood. Moreover, the absence of a predictive theory connecting catalysts to resulting chiralities has impeded systematic extension of this approach to other chiralities [8, 9, 10, 11, 12, 13].
Carbon nanotube growth proceeds via crystalline elongation from a nucleation cap—a curved graphene dome formed by incorporating six pentagons into a hexagonal lattice [14]. This cap architecture encodes complete chirality information within its topological structure [15, 16]. Developing a predictive theory connecting solid catalysts to the chiral outcomes requires addressing two interconnected challenges: deciphering chirality from cap topology, and elucidating how catalysts direct cap formation during nucleation. The first is a problem of topological interpretation, while the second involves complex physics of nucleation and catalyst-carbon interactions. Resolving these intertwined problems is essential for advancing this important field.
Here, we address the first challenge by introducing a topological theory that furnishes a mathematically rigorous connection between cap architecture and chirality—expressed as the vector sum rule. Building on this theoretical foundation, we then examine the second challenge through the lens of existing experimental evidence, proposing that chiral-selective growth on solid catalysts may be guided by structural matching between the carbon cap and catalyst surface. This integrated approach offers a coherent mechanistic framework for understanding the preferential formation of experimentally observed chiralities such as $(12,6)$ and $(14,4)$.
Figure 1: Schematic of a carbon cap structure. Red pentagons mark the positions of the pentagon carbon rings. Top left: the lattice basis vectors. Bottom left: the three topological basis vectors used in our framework.
Vector Sum Rule
To establish a rigorous connection between cap structure and nanotube chirality, we develop a topological framework that captures the essential geometry of carbon networks. A carbon nanotube is fundamentally a two-dimensional surface curved in three-dimensional space [17]. Within this framework, we define three basis topological vectors $\mathbf{e}_{i}$, $\mathbf{e}_{j}$, $\mathbf{e}_{k}$ (Fig. 1) and their respective inverses $\mathbf{e}_{i}^{-1}$, $\mathbf{e}_{j}^{-1}$, $\mathbf{e}_{k}^{-1}$. These vectors can be combined through a connection operation; for instance, $\mathbf{e}_{j}\cdot\mathbf{e}_{i}$ represents a basic building block, and any vector combined with its inverse yields the identity element $\mathbf{e}_{0}$. Repeated connections form topological paths, such as the closed path around a hexagonal ring:
$\overset{\leftrightarrow}{\mathbf{P}}_{6}=\mathbf{e}_{j}\cdot\mathbf{e}_{k}\cdot\mathbf{e}_{i}^{-1}\cdot\mathbf{e}_{j}^{-1}\cdot\mathbf{e}_{k}^{-1}\cdot\mathbf{e}_{i}.$
The set $S$, comprising the identity, all basis vectors, their inverses, and finite combinations under connection, contains every allowed topological path in graphene-like networks. To incorporate sublattice-mismatching defects, we define left and right rotation operators $\mathcal{L}$ and $\mathcal{R}$ that act on paths to introduce dislocations. These operators act cyclically on the basis vectors:
$\mathcal{L}$ $:\mathbf{e}_{j}=\mathcal{L}(\mathbf{e}_{i}),\quad\mathbf{e}_{k}=\mathcal{L}(\mathbf{e}_{j}),\quad\mathbf{e}_{i}^{-1}=\mathcal{L}(\mathbf{e}_{k})$
$\mathcal{R}$ $:\mathbf{e}_{k}^{-1}=\mathcal{R}(\mathbf{e}_{i}),\quad\mathbf{e}_{i}=\mathcal{R}(\mathbf{e}_{j}),\quad\mathbf{e}_{j}=\mathcal{R}(\mathbf{e}_{k})$
Although $S$ itself does not form a group, the subset generated from $\mathbf{e}_{0}$, $\mathbf{p}_{ji}=\mathbf{e}_{i}\cdot\mathbf{e}_{j}$, and $\mathbf{p}_{jk}=\mathbf{e}_{j}\cdot\mathbf{e}_{k}$ constitutes an Abelian group $V=\{\mathbf{e}_{0},\mathbf{p}_{ji},\mathbf{p}_{jk},\ldots\}^{+}$, wherein the connection operation is commutative and invertible. Within $V$, we adopt the plus sign to denote the connection operation between elements.
In graphene or nanotubes, any topological path connecting crystallographically equivalent atoms ($A\rightarrow A$ or $B\rightarrow B$) belongs to the group $V$. The value of such a path is defined as the element $\alpha\mathbf{p}_{ji}+\beta\mathbf{p}_{jk}$ (with $\alpha,\beta\in\mathbb{Z}$) to which it is homotopic. This leads to a purely topological definition of chirality: for a carbon nanotube of chirality $(n,m)$, a closed path encircling the circumference takes the value $n\mathbf{p}_{ji}+m\mathbf{p}_{jk}$ (see Supplemental Material S1 for detailed description of the topological framework).
We now incorporate pentagons to transform a closed graphene path into a nanotube path. Each pentagon $t$ ($t=1,\ldots,6$) occupies a lattice position $(n_{t},m_{t})$, represented topologically as $n_{t}\mathbf{p}_{ji}+m_{t}\mathbf{p}_{jk}$. The path from pentagon $t$ to pentagon $t+1$ (with cyclic indexing, $t+1\equiv 1$ for $t=6$) is given by:
$\mathbf{P}_{t\to t+1}=(n_{t+1}-n_{t})\mathbf{p}_{ji}+(m_{t+1}-m_{t})\mathbf{p}_{jk},$
Together, these form a closed graphene path summing to $\mathbf{e}_{0}$:
$\overset{\leftrightarrow}{\mathbf{P}}_{\text{Graphene}}=\sum_{t=1}^{6}\mathbf{P}_{t\to t+1}=\mathbf{e}_{0}.$
Introducing a pentagon—a topological kink represented by the path
$\mathbf{P}_{5}=\mathbf{e}_{j}\cdot\mathbf{e}_{k}\cdot\mathbf{e}_{i}^{-1}\cdot\mathbf{e}_{j}^{-1}\cdot\mathbf{e}_{k}^{-1}=\mathbf{e}_{i}^{-1}$
—breaks this closure. Although $\mathbf{P}_{1\to 2}$ connects directly to $\mathbf{P}_{5}$, the combined path $\mathbf{P}_{1\to 2}+\mathbf{P}_{5}$ cannot connect directly to $\mathbf{P}_{2\to 3}$ due to a fundamental sublattice mismatch introduced by the pentagonal defect. This topological obstruction is resolved by systematically applying right rotations $\mathcal{R}$ to subsequent segments, which restores sublattice registry. The complete circumferential path of the nanotube, incorporating all six pentagons, is therefore:
$\overset{\leftrightarrow}{\mathbf{P}}_{\text{CNT}}=\sum_{t=1}^{6}\mathcal{R}^{t-1}(\mathbf{P}_{t\to t+1}+\mathbf{P}_{5}),$ (1)
where $\mathcal{R}^{t-1}$ denotes t1t-1 successive applications of the $\mathcal{R}$ operator.
Given that three right rotations invert any path, $\mathcal{R}^{3}(\mathbf{P})=\mathbf{P}^{-1}$, the pentagon contributions sum to identity: $\sum_{t=1}^{6}\mathcal{R}^{t-1}(\mathbf{P}_{5})=\mathbf{e}_{0}$. Substituting the explicit path forms yields:
$\overset{\leftrightarrow}{\mathbf{P}}_{\text{CNT}}=\sum_{t=1}^{6}\mathcal{R}^{t-1}(n_{t}\mathbf{p}_{ji}+m_{t}\mathbf{p}_{jk}).$ (2)
To simplify the expression in Equation (2), we exploit the identity $\mathbf{P}=\mathcal{R}^{t-1}[\mathcal{L}^{t-1}(\mathbf{P})]$ and re-parameterize each $(n_{t},m_{t})$ in successively left-rotated topological spaces. This aligns all pentagon contributions into a common reference frame:
$n_{t}\mathbf{p}_{ji}+m_{t}\mathbf{p}_{jk}=\dot{n}_{t}\mathcal{L}^{t-1}(\mathbf{p}_{ji})+\dot{m}_{t}\mathcal{L}^{t-1}(\mathbf{p}_{jk}).$ (3)
The closed nanotube path defines its chirality:
$\overset{\leftrightarrow}{\mathbf{P}}_{\text{CNT}}=n\cdot\mathbf{p}_{ji}+m\cdot\mathbf{p}_{jk}.$ (4)
Substituting Equations (3) and (4) into Equation (2) gives:
$n\cdot\mathbf{p}_{ji}+m\cdot\mathbf{p}_{jk}=\left(\sum_{t=1}^{6}\dot{n}_{t}\right)\mathbf{p}_{ji}+\left(\sum_{t=1}^{6}\dot{m}_{t}\right)\mathbf{p}_{jk}.$ (5)
Figure 2: Chiral index assignment of a carbon nanotube cap. (Top left) Six rotationally generated vector spaces obtained by successive left rotations. (Bottom) Chiral indices of the cap mapped onto the graphene lattice in corresponding vector spaces. (Top right) Atomic configuration of the cap and the resulting carbon nanotubes, with the cap-nanotube boundary defined by the presence of a stable edge structure during nucleation (see Supplemental Material S2 for details)
Equation (5) reveals the fundamental linear relationship. The Abelian group $V$ is isomorphic to the group generated by physical lattice vectors $\mathbf{r}_{n}$, $\mathbf{r}_{m}$. Here, $\mathbf{p}_{ji}$ and $\mathbf{p}_{jk}$ are their topological proxies; a left rotation $\mathcal{L}$ corresponds to a $\pi/3$ real-space rotation of the lattice basis. We therefore introduce a chiral index notation for the cap: $(\dot{n}_{1},\dot{m}_{1}),(\dot{n}_{2},\dot{m}_{2}),\ldots,(\dot{n}_{6},\dot{m}_{6})]$, where each tuple $(\dot{n}_{t},\dot{m}_{t})$ represents the coordinates of a pentagon within the lattice space $(\mathbf{r}_{n}^{t},\mathbf{r}_{m}^{t})$ (Fig. 2). Treating these notation tuples as integer vectors leads to the final vector sum rule:
$(n,m)=\sum_{t=1}^{6}\left(\dot{n}_{t},\dot{m}_{t}\right).$ (6)
Equation (6) presents a central result of this work: the vector sum rule governing carbon nanotube chirality. This rule dictates that the chiral indices $(n,m)$ are determined entirely by the vector sum of the coordinates of the six pentagons, evaluated across rotationally related lattice vector spaces. A profound implication is that while a given cap can be represented by multiple notation sequences due to the symmetry of the hexagonal lattice, the resulting sum in Eq. (6) remains invariant under all symmetry operations of the P6/mmm space group. This means that although the specific notation sequences of a cap may be altered by symmetry operations—including translation, rotation, and reflection—the topological sum rule itself remains unchanged. This invariance underscores the rule’s fundamental character as a universal topological law.
The vector sum rule provides not only an intuitive bridge between cap structure and nanotube chirality but also profound physical insight into its origin. We propose a topological interpretation in which the graphene edge is taken as a reference state, since any closed path on its lattice is homotopic to the identity element $\mathbf{e}_{0}$. Each pentagon, however, acts as a topological deflector, rotating the growing edge path by $\pi/3$—a direct consequence of the sublattice mismatch introduced by the pentagonal defect. After six discrete rotations, the crystalline registry is restored but the edge is redirected to form a nanotube circumference. The chirality of the resulting nanotube is determined by the spatial distribution of these six discrete rotations within the hexagonal lattice.
Due to the translational symmetry of the hexagonal lattice, fully defining a cap requires exactly 10 integers when one pentagon is taken as a reference position. Interestingly, a cap contains significantly more structural information than its resulting nanotube edge, which is completely described by only two integers $(n,m)$. As a consequence, each chirality corresponds to a large number of distinct cap configurations. For example, $(12,6)$ nanotubes can theoretically nucleate from 22,127 different caps [8], revealing substantial structural degeneracy in the nucleation process.
Deterministic Nucleation
Figure 3: (a)-(c) Schematic illustrations of lattice matching between nanotube caps and $\langle 1,1,1\rangle$ crystallographic facets of catalyst nanoparticles. Cap notations $\ldots], [(2,1),(2,1),$(3,0),(2,1),(2,1),(3,0)][(2,1),(2,1),(3,0),(2,1),(2,1),(3,0)], and $(3,0),\ldots]$ correspond to $(12,6)$, $(14,4)$, and $(18,0)$ nanotubes, respectively. (d) Per-carbon-atom formation energy of caps on copper catalyst surface as a function of chiral angle $\theta$, determined by molecular dynamics simulations enumerating 730 distinct configurations corresponding to 18 chiralities. The initial graphene lattice contains 19 hexagons, with pentagon incorporation occurring externally at the edge. Atomic interactions are described by a bond-order potential [18], under NVT ensemble with temperature maintained at 1 K throughout all simulations. The color bar shows the nucleation probabilities derived from a simplified model.
Two distinct pathways—clone growth and solid catalyst–specified growth via CVD—offer routes toward atomic-level control over nanotube synthesis. Clone growth directly programs the cap architecture through molecular-scale design. For instance, chemical opening of $\mathrm{C}_{60}$ molecules yields well-defined cap sequences such as $(1,0),(2,0),\ldots]$ and $(0,0),(1,1),\ldots]$, which template the growth of $(9,0)$ and $(5,5)$ nanotubes, respectively (Fig. 2[19]. Alternatively, molecular templates synthesized through organic chemistry—such as the $\mathrm{C}_{96}\mathrm{H}_{54}$ molecule, which preorganizes the sequence $(2,1),(0,1),\ldots]$—can direct cap assembly to selectively form $(6,6)$ nanotubes [20]. Clone growth provides a clear demonstration of the direct pathway from nucleation to nanotube formation. However, it suffers from slow elongation rates beyond the microscale.
In contrast, CVD using solid catalysts enables rapid elongation, but nucleation remains poorly understood, despite having produced several nanotubes with high enantiomeric purity [21, 22]. Among these high-purity species, the $(12,6)$ species exhibits notable and reproducible abundance across a wide range of synthesis conditions [23, 6, 24]. Although kinetic theory accounts for its rapid elongation, the preferential nucleation of $(12,6)$ nanotubes has remained an experimental puzzle. Here, we address this long-standing question through a topological framework.
First, we need to establish the concept of deterministic nucleation by addressing a fundamental question: is chirality a statistical outcome of stochastic nucleation processes, or is it deterministically determined by a specific nucleation process that directs growth toward a preferred $(n,m)$? This question is subtle yet critically important. The unconscious assumption of the former has led to an overwhelming focus on post-nucleation explanations of chirality control, particularly through edge-matching mechanisms. This issue relates to the fact that each chirality corresponds to a large number of distinct cap configurations, with this number growing rapidly with diameter. Eq. (6) establishes that any change in pentagon position induces a deterministic and quantifiable shift in chirality, unless compensated by countervailing changes elsewhere in the cap. For instance, shifting one pentagon from $(0,0)$ to $(1,1)$ in a $(5,5)$ nanotube redirects growth to a $(6,6)$ nanotube (Fig. 2). Restoring the original $(5,5)$ chirality would require compensating changes in other pentagon positions summing to (1,1)(-1,-1)—a scenario of low statistical probability in an entropy-dominated regime. This implies that if the nucleation process equally generated all possible caps for subsequent growth, it would also equally generate caps for neighboring chiralities through minor pentagon position shifts. Such behavior is never observed experimentally. Chirality enrichment follows specific patterns that cannot be explained by proximity in chirality space. For example, while $(12,6)$ nanotubes are enriched, $(12,7)$ nanotubes have not been observed enriching in any experimental studies. From a broader perspective, the total number of caps meeting the diameter requirements proximate to $(12,6)$ nanotubes reaches millions, yet only a very small percentage correspond specifically to $(12,6)$. Statistically, it is impossible for a random nucleation process to result in such chirality enrichment. The origin of chirality control must therefore stem from deterministic nucleation where only a few (or one) closely related caps are engineered—as unambiguously evidenced by the precise structural templating observed in clone growth experiments [19, 20].
The remaining challenges are to identify the specific cap architectures responsible for the preferential formation of $(12,6)$ nanotubes and to understand the underlying physical principles that favor such configurations. The vector sum rule provides the foundation to systematically enumerate all possible caps satisfying this chirality condition under catalyst constraints, enabling targeted investigation of their relative stabilities and catalytic matching properties..
carbon caps exhibiting six-fold rotational symmetry. These caps correspond to nanotubes with chiral indices $(6n,6m)$ and are topologically represented by the notation $(n,m),\ldots]$, as exemplified by $(1,0),\ldots]$, $(1,1),\ldots]$, and $(2,1),\ldots]$. Such symmetric caps achieve coherent structural registry with $\langle 1,1,1\rangle$ crystallographic facets of various face-centered catalysts, with all six pentagons positioned along the cap periphery (Fig. 3a). This configuration creates favorable nucleation conditions, enabling epitaxial extension of the cap across the catalyst surface in a graphene-like growth mode [25, 26]. Pentagon incorporation is triggered either when the expanding carbon network encounters the catalyst boundary or when progressive lattice mismatch during epitaxial growth induces large-scale structural strain [27, 28, 29, 30].
The enrichment of $(12,6)$ nanotubes can thus be explained by the combination of optimal nucleation characteristics with favorable post-nucleation growth kinetics—specifically, a diameter of approximately 1.3 nm and a chiral angle of $19.1^{\circ}$ [31, 32, 33].
However, achieving such well-defined growth requires precise incorporation of all six pentagons at $(2,1)$ sublattice sites (or equivalently, $(1,2)$ sites, which yield the mirror-symmetric $(6,12)$ chirality). Any deviation from this ideal configuration—for example, substituting two $(2,1)$ pentagons with $(3,0)$ sites—deterministically shifts the resulting chirality to the semiconducting $(14,4)$ type (Fig. 3b). Consistent with this theoretical framework, the $(14,4)$ nanotube exhibits experimental enrichment under similar synthesis conditions [7]. Moreover, several other structurally related chiralities—including $(18,0)$, $(16,2)$, $(15,3)$, and (9,9)(9,9)—are also enriched under specific synthesis conditions [29].
Yet a deeper question persists: why are $(12,6)$ nanotubes overwhelmingly preferred over other chiralities across different synthesis conditions? Although a rapid post-nucleation elongation rate makes them easier to detect experimentally, this alone cannot explain their abundance across diverse synthesis conditions unless preferential nucleation conditions also exist relative to neighboring caps. Among the 730 distinct cap configurations (satisfying the Isolated Pentagon Rule) closely related to this six-fold symmetry, only 116 configurations yield $(12,6)$ chirality. From a configurational entropy perspective alone, other chiralities should be substantially more probable within this symmetry family. The experimental predominance of $(12,6)$ nanotubes therefore points to decisive energetic advantages and kinetic privilege of the $(2,1),\ldots]$ cap that override entropic factors.
As revealed by in situ transmission electron microscopy [26], during nucleation, edges undergo repeated etching and regrowth, gradually adjusting toward lower energy states. In this stage, low-energy cap configurations gain significant advantage over alternatives. As shown in Fig. 3c, the formation energy of caps registered with crystallographic facets approximately follows a scaled sinusoidal function of the chiral angle, with the $(2,1),\ldots]$ cap exhibiting the largest deviation from this trend. This particular advantage of the $(2,1),\ldots]$ cap arises from two key factors: maximized inter-pentagon distances that reduce strain energy, and closer registry with catalyst surfaces that minimizes interfacial energy. Furthermore, nucleation is not entirely an equilibrium process; the kinetics of pentagon incorporation also play a crucial role. Experimental evidence indicates that incorporating a pentagon by adding a $\text{C}_{3}$ atom sequence to an anti-armchair site is significantly less favorable than adding a $\text{C}_{2}$ atom to a zigzag site [30]. Consequently, the incorporation of a pentagon at a $(3,0)$ site requires either its right neighboring $(1,2)$ site or left neighboring $(2,1)$ site to incorporate a hexagon first, thereby converting the $(3,0)$ site to a zigzag configuration.
To quantify this kinetic barrier, we consider a simplified model where the probability of pentagon incorporation at a zigzag site is $\phi$ ($0<\phi<1$), while the probability of hexagon incorporation at the same site is $1-\phi$. The overall probability of pentagon incorporation at a $(3,0)$ site then becomes $2(1-\phi)\phi$. Because incorporating a hexagon at a zigzag site also requires a $\text{C}_{3}$ atom sequence, which is less favorable than the $\text{C}_{2}$ addition for pentagon incorporation, we have $1-\phi<\phi$. Therefore, the probability $2(1-\phi)\phi$ is lower than the probability $\phi$ for direct pentagon incorporation at $(2,1)$ or $(1,2)$ sites. The total probability to form a cap, ignoring probability fluctuations from other incorporation events, is given by $\Psi=\prod_{t=1}^{6}\phi_{t}$, where $\phi_{t}$ represents the incorporation probability at each of the six pentagon sites. $\Psi$ represents the relative formation probability of a specific cap compared to other configurations. This simple kinetic analysis confirms that the $(2,1),\ldots]$ cap is privileged and favored, while pro-zigzag caps are significantly less favorable to nucleate despite having lower formation energy in some cases, as illustrated in Fig. 3c (assuming $\phi=0.9$ for demonstration purposes). From a stochastic perspective, although chirality is deterministically encoded by each cap, the emergence of a distribution of chiralities during synthesis is statistically inevitable, given that the probability to form unfavored caps is low but non-zero. Of course, an accurate model needs to consider all incorporation pathways and, most importantly, account for how morphological changes of the catalyst surface influence the incorporation probabilities. This critical issue has been a main focus in graphene synthesis toward large-scale pristine lattice, being studied by in situ transmission electron microscopy [30]. In graphene studies, pentagons and heptagons typically emerge as pairs, whereas in nanotubes, the nucleation process involves exclusively pentagons. This remains an uncharted territory; therefore, a concise study of pentagon incorporation dynamics is required [34, 35].
Beyond nucleation control, our topological framework opens pathways for the rational design of carbon-based heterostructures, including intramolecular nanotube junctions and graphene-nanotube hybrids (see Supplemental Material S3). By strategically incorporating pentagonal and heptagonal defects into the hexagonal lattice, diverse target architectures can be systematically constructed. Thus, our theory serves as a versatile and predictive tool for the atomic-scale engineering of complex carbon nanomaterials.
Conclusion
Through the discovery of the vector sum rule and its application to six-fold symmetric nucleation caps, we have reframed chirality-controlled nanotube synthesis as a deterministic process governed by cap topology and epitaxial matching with catalyst surfaces. While previous theories attributed chirality-selective growth primarily to edge matching mechanisms [36], our findings establish that nucleation is deterministically controlled by these topological and interfacial factors. This theoretical framework coherently explains the experimental enrichment of $(12,6)$ nanotubes while accounting for the emergence of chiral enrichment in neighboring species. The parallel enrichment observed around $(6,6)$ nanotubes—which similarly feature six-fold symmetric caps—further validates this mechanism [37]. Although our work captures the essential features of nucleation, developing predictive synthesis requires a more comprehensive understanding of how catalyst morphology directs pentagon formation. Our results provide the theoretical foundation for advancing toward this goal.
Acknowledgements
Fundings: This study was supported by the National Natural Science Foundation of China (Nos. 12462009 and 12132008). We thank Nader Anani for advice on the manuscript.
Data avalibility: All data are available in the main text or the supplementary materials. A direct implementation of the topological framework into computer code for designing carbon caps is available at GitHub (https://github.com/ZhengrongGuoChina/Graphene-topology).